# Contraherent cosheaves of contramodules on Noetherian formal schemes

### 29 Mar 2026 v1

@@ 3.4. Quasi-coherent torsion sheaves
We will call condition (i) is the torsion axiom, while condition (ii) will be called the quasi-coherence axiom. This definition is to be compared with the Enochs–Estrada description of quasi-coherent sheaves on schemes [6, Section 2], [29, Section 2.2].
@@
Recall that $\mathfrak{O}(\mathfrak{V})$ is a flat contramodule $\mathfrak{O}(\mathfrak{U})$-module by Lemma 2.8.6(3) (and a flat $\mathfrak{O}(\mathfrak{U})$-module by Lemma 2.5.10(6) or (7)). Hence the tensor product functor $\mathfrak{O}(\mathfrak{V})\otimes_{\mathfrak{O}(\mathfrak{U})}{-}$ preserves exactness of the left exact sequence (15), and we obtain a left exact sequence of $\mathfrak{O}(\mathfrak{V})$-modules
@@
Let $\mathcal{M}$ be a quasi-coherent torsion sheaf on $\mathbf{B}$, let $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ be affine open subschemes, and let $\mathfrak{W}\subset\mathfrak{U}$ be an affine open subscheme subordinate to $\mathbf{W}$. Then the restriction map of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})$-modules $\mathcal{M}(\mathfrak{W})\longrightarrow\mathcal{M}(\mathfrak{V}\cap\mathfrak{W})$ induces an isomorphism of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})$-modules $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathcal{M}(\mathfrak{W})\longrightarrow\mathcal{M}(\mathfrak{V}\cap\mathfrak{W})$.
@@
is an isomorphism. Now the map (18) becomes an isomorphism after the adic completion functor $\Lambda$ is applied, while the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})$-module $\mathcal{M}(\mathfrak{W})$ is torsion, and it remains to refer to Corollary 2.5.4. ∎
Finally, let $\mathcal{M}$ be a sheaf of torsion $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathfrak{X}$. We will say that $\mathcal{M}$ is a quasi-coherent torsion sheaf on $\mathfrak{X}$ if the quasi-coherence axiom (ii) holds for all affine open subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ (or equivalenlty, for all affine open subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$). According to the preceding discussion in this section, the category of quasi-coherent torsion sheaves on $\mathfrak{X}$ is equivalent to the category of quasi-coherent torsion (pre)sheaves on the topology base $\mathbf{B}$ consisting of all affine open formal subschemes of $\mathfrak{X}$ subordinate to $\mathbf{W}$ (for any given open covering $\mathbf{W}$ of the formal scheme $\mathfrak{X}$).
@@
A short sequence $0\longrightarrow\mathcal{L}\longrightarrow\mathcal{M}\longrightarrow\mathcal{N}\longrightarrow 0$ of quasi-coherent torsion sheaves on $\mathfrak{X}$ is exact in $\mathfrak{X}{\operatorname{\mathsf{--Tors}}}$ if and only if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the short sequence of torsion $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $0\longrightarrow\mathcal{L}(\mathfrak{U})\longrightarrow\mathcal{M}(\mathfrak{U})\longrightarrow\mathcal{N}(\mathfrak{U})\longrightarrow 0$ is exact. By Lemma 2.13.8(a), it suffices to check this condition for affine open subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$.
Moreover, a complex $\mathcal{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathfrak{X}{\operatorname{\mathsf{--Tors}}}$ is exact if and only if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the complex of torsion $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules is exact in $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U}){\operatorname{\mathsf{--Tors}}}$. By Lemma 2.13.8(d), it suffices to check this condition for affine open subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$.
@@ 3.5. Contraherent cosheaves of contramodules
We keep the notation of Section 3.4. So $\mathfrak{X}$ is a locally Noetherian formal scheme, $\mathbf{W}$ is an open covering of $\mathfrak{X}$, and $\mathbf{B}$ is the base of open subsets in $\mathfrak{X}$ consisting of all the affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$.
@@
We will call condition (iii) is the contramoduleness axiom, while condition (iv) will be called the contraherence axiom, and condition (v) will be called the contraadjustedness axiom. Cf. the definitions of a (locally) cohereherent cosheaf of modules over a scheme in [17, Sections 2.2 and 3.1] and [29, Section 2.2].
@@
The notion of a contraherent cosheaf over a scheme (hence also of a contraherent cosheaf of contramodules over a formal scheme) is not local. See [17, Example 3.2.1] for a counterexample.
More precisely, in the context of a formal scheme $\mathfrak{X}$, for any contraherent cosheaf of contramodules $\mathfrak{P}$ on $\mathbf{B}$ extended to a cosheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathfrak{P}$ on the whole of $\mathfrak{X}$, both the contramoduleness axiom (iii) and the contraadjustedness axiom (v) hold for all affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$. For the contramoduleness axiom, this was mentioned in the previous paragraph. For the contraadjustedness axiom, this follows from the facts that the restrictions of scalars preserve contraadjustedness (see Lemma 2.9.7), finite direct sums of contraadjusted contramodules are contraadjusted, and any quotient contramodule of a contraadjusted contramodule is contraadjusted (as mentioned in Section 2.9). It is the contraherence axiom (iv) that is not local.
@@ 3.6. Locally contraherent cosheaves of contramodules
Let $\mathfrak{P}$ be a cosheaf of contramodule $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathfrak{X}$. We will say that $\mathfrak{P}$ is a $\mathbf{W}$-locally contraherent cosheaf of contramodules (or a $\mathbf{W}$-locally contraherent contramodule cosheaf) on $\mathfrak{X}$ if the contraherence axiom (iv) and the contraadjustedness axiom (v) hold for all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$. If this is the case, then the contraadjustedness axiom actually holds for all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$, as explained in Section 3.5. According to the discussion in Section 3.5, the category of $\mathbf{W}$-locally contraherent cosheaves of contramodules on $\mathfrak{X}$ is equivalent to the category of contraherent co(pre)sheaves on the topology base $\mathbf{B}$ consisting of all affine open subschemes of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
@@
The next proposition is the formal scheme version of the homological criterion of contraherence from [17, Lemma 3.2.2]. The definition of the homological Čech complex $C_{\text{\smaller\smaller$\scriptstyle\bullet$}}(\{\mathfrak{V}_{\alpha}\},\mathfrak{P})$ was given in Section 3.2.
@@
is an isomorphism. Finally, the map (23) becomes an isomorphism after the adic completion functor $\Lambda$ is applied, while the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$-module $\mathfrak{P}[\mathfrak{W}]$ is a (quotseparated) contramodule, and it remains to refer to Corollary 2.5.7 (whose flatness assumption is obviously satisfied for the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$-module $F=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$). ∎
@@
A short sequence of $\mathbf{W}$-locally contraherent cosheaves of contramodules $0\longrightarrow\mathfrak{L}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ on $\mathfrak{X}$ is said to be (admissible) exact if the short sequence of (contramodule) $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $0\longrightarrow\mathfrak{L}[\mathfrak{U}]\longrightarrow\mathfrak{M}[\mathfrak{U}]\longrightarrow\mathfrak{N}[\mathfrak{U}]\longrightarrow 0$ is exact for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$. By Lemma 2.13.10(a), it suffices to check this condition for affine open subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
@@
Furthermore, it follows from Lemma 2.13.10(b) that the full subcategory $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}$ is closed under kernels of admissible epimorphisms in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. More generally, the full subcategory $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is closed under kernels of admissible epimorphisms in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}$. So a morphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is an admissible epimorphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ if and only if it is an admissible epimorphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}$. The similar assersions for admissible monomorphisms are not true [17, Example 3.2.1].
@@ 3.7. Locally cotorsion locally contraherent cosheaves
A $\mathbf{W}$-locally contraherent cosheaf of contramodules $\mathfrak{P}$ on $\mathfrak{X}$ is said to be locally cotorsion if the contramodule $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathfrak{P}[\mathfrak{U}]$ is cotorsion for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$. By Proposition 2.12.4 and Corollary 2.12.12, it suffices to check that condition for affine open subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
We will denote the full subcategory of locally cotorsion $\mathbf{W}$-locally contraherent cosheaves of contramodules on $\mathfrak{X}$ by $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}\subset\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. The category of locally cotorsion contraherent cosheaves of contramodules on $\mathfrak{X}$ will be denoted by $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}^{\mathsf{lct}}=\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\{\mathfrak{X}\}}^{\mathsf{lct}}$. It follows from the previous paragraph that $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}^{\mathsf{lct}}=\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}\cap\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$ and $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}=\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}\cap\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}^{\mathsf{lct}}$. The category of locally cotorsion locally contraherent cosheaves of contramodules on $\mathfrak{X}$ will be denoted by $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}^{\mathsf{lct}}=\bigcup_{\mathbf{W}}\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$.
@@
(a) A complex $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in the exact category of $\mathbf{W}$-locally contraherent cosheaves of contramodules $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is exact if and only if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the complex of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}[\mathfrak{U}]$ is exact. It suffices to check this condition for affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ belonging to any chosen open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
(a) A complex $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in the exact category of locally cotorsion $\mathbf{W}$-locally contraherent cosheaves of contramodules $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$ is exact if and only if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the complex of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}[\mathfrak{U}]$ is exact. It suffices to check this condition for affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ belonging to any chosen open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.

### 17 Apr 2026 v2

@@ 3.4. Quasi-coherent torsion sheaves
We will call condition (i) is the torsion axiom, while condition (ii) will be called the quasi-coherence axiom. This definition is to be compared with the Enochs–Estrada description of quasi-coherent sheaves on schemes [7, Section 2], [30, Section 2.2].
@@
Recall that $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})$ is a flat contramodule $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module by Lemma 2.8.6(3) (and a flat $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module by Lemma 2.5.10(6) or (7)). Hence the tensor product functor $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}{-}$ preserves exactness of the left exact sequence (15), and we obtain a left exact sequence of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})$-modules
@@
Let $\mathcal{M}$ be a quasi-coherent torsion sheaf on $\mathbf{B}$, let $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ be affine open formal subschemes, and let $\mathfrak{W}\subset\mathfrak{U}$ be an affine open formal subscheme subordinate to $\mathbf{W}$. Then the restriction map of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})$-modules $\mathcal{M}(\mathfrak{W})\longrightarrow\mathcal{M}(\mathfrak{V}\cap\mathfrak{W})$ induces an isomorphism of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})$-modules $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathcal{M}(\mathfrak{W})\longrightarrow\mathcal{M}(\mathfrak{V}\cap\mathfrak{W})$.
@@
is an isomorphism. Now the map (18) becomes an isomorphism after the adic completion functor $\Lambda$ is applied, while the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})$-module $\mathcal{M}(\mathfrak{W})$ is torsion, and we can refer to Corollary 2.5.4. ∎
Finally, let $\mathcal{M}$ be a sheaf of torsion $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathfrak{X}$. We will say that $\mathcal{M}$ is a quasi-coherent torsion sheaf on $\mathfrak{X}$ if the quasi-coherence axiom (ii) holds for all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ (or equivalently, for all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$). According to the preceding discussion in this section, the category of quasi-coherent torsion sheaves on $\mathfrak{X}$ is equivalent to the category of quasi-coherent torsion (pre)sheaves on the topology base $\mathbf{B}$ consisting of all affine open formal subschemes of $\mathfrak{X}$ subordinate to $\mathbf{W}$ (for any given open covering $\mathbf{W}$ of the formal scheme $\mathfrak{X}$).
@@
A short sequence $0\longrightarrow\mathcal{L}\longrightarrow\mathcal{M}\longrightarrow\mathcal{N}\longrightarrow 0$ of quasi-coherent torsion sheaves on $\mathfrak{X}$ is exact in $\mathfrak{X}{\operatorname{\mathsf{--Tors}}}$ if and only if, for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$, the short sequence of torsion $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $0\longrightarrow\mathcal{L}(\mathfrak{U})\longrightarrow\mathcal{M}(\mathfrak{U})\longrightarrow\mathcal{N}(\mathfrak{U})\longrightarrow 0$ is exact. By Lemma 2.17.8(a), it suffices to check this condition for affine open formal subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$.
Moreover, a complex $\mathcal{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathfrak{X}{\operatorname{\mathsf{--Tors}}}$ is exact if and only if, for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$, the complex of torsion $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules is exact in $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U}){\operatorname{\mathsf{--Tors}}}$. By Lemma 2.17.8(d), it suffices to check this condition for affine open formal subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$.
@@ 3.5. Contraherent cosheaves of contramodules
We keep the notation of Section 3.4. So $\mathfrak{X}$ is a locally Noetherian formal scheme, $\mathbf{W}$ is an open covering of $\mathfrak{X}$, and $\mathbf{B}$ is the base of open subsets in $\mathfrak{X}$ consisting of all the affine open formal subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$.
@@
We will call condition (iii) is the contramoduleness axiom, while condition (iv) will be called the contraherence axiom, and condition (v) will be called the contraadjustedness axiom. Cf. the definitions of a (locally) cohereherent cosheaf of modules over a scheme in [18, Sections 2.2 and 3.1] and [30, Section 2.2].
@@
The notion of a contraherent cosheaf over a scheme (hence also of a contraherent cosheaf of contramodules over a formal scheme) is not local. See [18, Example 3.2.1] for a counterexample.
More precisely, in the context of a formal scheme $\mathfrak{X}$, for any contraherent cosheaf of contramodules $\mathfrak{P}$ on $\mathbf{B}$ extended to a cosheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathfrak{P}$ on the whole of $\mathfrak{X}$, both the contramoduleness axiom (iii) and the contraadjustedness axiom (v) hold for all affine open formal subschemes $\mathfrak{U}\subset\mathfrak{X}$. For the contramoduleness axiom, this was mentioned in the previous paragraph. For the contraadjustedness axiom, this follows from the facts that the restrictions of scalars preserve contraadjustedness (see Lemma 2.12.7), finite direct sums of contraadjusted contramodules are contraadjusted, and any quotient contramodule of a contraadjusted contramodule is contraadjusted (as mentioned in Section 2.12). It is the contraherence axiom (iv) that is not local.
@@ 3.6. Locally contraherent cosheaves of contramodules
Let $\mathfrak{P}$ be a cosheaf of contramodule $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathfrak{X}$. We will say that $\mathfrak{P}$ is a $\mathbf{W}$-locally contraherent cosheaf of contramodules (or a $\mathbf{W}$-locally contraherent contramodule cosheaf) on $\mathfrak{X}$ if the contraherence axiom (iv) and the contraadjustedness axiom (v) hold for all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$. If this is the case, then the contraadjustedness axiom actually holds for all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$, as explained in Section 3.5. According to the discussion in Section 3.5, the category of $\mathbf{W}$-locally contraherent cosheaves of contramodules on $\mathfrak{X}$ is equivalent to the category of contraherent co(pre)sheaves on the topology base $\mathbf{B}$ consisting of all affine open formal subschemes of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
@@
The next proposition is the formal scheme version of the homological criterion of contraherence from [18, Lemma 3.2.2]. The definition of the homological Čech complex $C_{\text{\smaller\smaller$\scriptstyle\bullet$}}(\{\mathfrak{V}_{\alpha}\},\mathfrak{P})$ was given in Section 3.2.
@@
is an isomorphism. Finally, the map (23) becomes an isomorphism after the adic completion functor $\Lambda$ is applied, while the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$-module $\mathfrak{P}[\mathfrak{W}]$ is a (quotseparated) contramodule, and we can refer to Corollary 2.5.7 (whose flatness assumption is obviously satisfied for the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$-module $F=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$). ∎
@@
A short sequence of $\mathbf{W}$-locally contraherent cosheaves of contramodules $0\longrightarrow\mathfrak{L}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ on $\mathfrak{X}$ is said to be (admissible) exact if the short sequence of (contramodule) $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $0\longrightarrow\mathfrak{L}[\mathfrak{U}]\longrightarrow\mathfrak{M}[\mathfrak{U}]\longrightarrow\mathfrak{N}[\mathfrak{U}]\longrightarrow 0$ is exact for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$. By Lemma 2.17.10(a), it suffices to check this condition for affine open formal subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
@@
Furthermore, it follows from Lemma 2.17.10(b) that the full subcategory $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}$ is closed under kernels of admissible epimorphisms in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. More generally, the full subcategory $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is closed under kernels of admissible epimorphisms in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}$. So a morphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is an admissible epimorphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ if and only if it is an admissible epimorphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}$. The similar assersions for admissible monomorphisms are not true [18, Example 3.2.1].
@@ 3.7. Locally cotorsion locally contraherent cosheaves
A $\mathbf{W}$-locally contraherent cosheaf of contramodules $\mathfrak{P}$ on $\mathfrak{X}$ is said to be locally cotorsion if the contramodule $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathfrak{P}[\mathfrak{U}]$ is cotorsion for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$. By Proposition 2.16.4 and Corollary 2.16.12, it suffices to check that condition for affine open formal subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
We will denote the full subcategory of locally cotorsion $\mathbf{W}$-locally contraherent cosheaves of contramodules on $\mathfrak{X}$ by $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}\subset\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. The category of locally cotorsion contraherent cosheaves of contramodules on $\mathfrak{X}$ will be denoted by $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}^{\mathsf{lct}}=\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\{\mathfrak{X}\}}^{\mathsf{lct}}$. It follows from the previous paragraph that $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}^{\mathsf{lct}}=\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}\cap\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$ and $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}=\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}\cap\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}^{\mathsf{lct}}$ (in the notation from the end of the previous section). The category of locally cotorsion locally contraherent cosheaves of contramodules on $\mathfrak{X}$ will be denoted by $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}^{\mathsf{lct}}=\bigcup_{\mathbf{W}}\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$.
@@
(a) A complex $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in the exact category of $\mathbf{W}$-locally contraherent cosheaves of contramodules $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is exact if and only if, for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the complex of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}[\mathfrak{U}]$ is exact. It suffices to check this condition for affine open formal subschemes $\mathfrak{U}\subset\mathfrak{X}$ belonging to any chosen open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
(a) A complex $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in the exact category of locally cotorsion $\mathbf{W}$-locally contraherent cosheaves of contramodules $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$ is exact if and only if, for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the complex of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}[\mathfrak{U}]$ is exact. It suffices to check this condition for affine open formal subschemes $\mathfrak{U}\subset\mathfrak{X}$ belonging to any chosen open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.